OpenStudy (anonymous):

multiply and simplify x^2-4/x^2-1*x+1/x+2

OpenStudy (ivettef365):

so you have x^2-4 x+1 ------ x ------- x^2-1 x+2

OpenStudy (anonymous):

yeahh

OpenStudy (ivettef365):

ok x^2-4 = (a^2-b^2) = (a+b)(a-b), so you have (x^2-2^2) (x+2)(x-2) x+1 --------- x ------ (x-1)(x+1) x+2

OpenStudy (ivettef365):

now you can cancel top and bottom the (x+1) and also the (x+2) what is left?

OpenStudy (anonymous):

(x+2(x-2) --------- (x-1)(x+1)

OpenStudy (anonymous):

? @ivettef365

OpenStudy (ivettef365):

x+2 cancels so you have on top x-2 x+1 cancels so you have on bottom x-1 so the answer would be x - 2 ----- x -1

OpenStudy (anonymous):

ohhhh okay , thanks!!

OpenStudy (ivettef365):

yw :)